OpenStudy (anonymous):

solve: (log (base 3) 5x)-(log (base 3) 1+sqrt(x))=2

OpenStudy (anonymous):

\[log_35x-log_31+\sqrt{x}=2\] \[log_3\frac{5x}{1}+\sqrt{x}=2\] \[log_35x+\sqrt{x}=2\] is it log(sqrt(x) or log(sqrt(x))?

OpenStudy (anonymous):

Its Log base 3 (1+sqrt(x))